A cukaryotic pre-mRNA transcript has 5 exoms and 4
Answer #1

During post transcriptional modification of eukaryotic pre-mRNA introns are removed and exons are rearranged in a specific order to generate a matured mRNA with required sequence by alternate splicing.

Number of nucleotides in mRNA = Sum of nucleotide length in (Exon 1 + Exon 2 + Exon 3 + Exon 4)

Number of nucleotides in mRNA = 219 + 54 + 162 + 45

Number of nucleotides in mRNA = 480

Since three nucleotides collectively (triplet codon) code for a single amino acid number of amino acids in the given protein = Number of nucleotides / 3 = 480 / 3 = 160 amino acids.

Molecular weight of one amino acid is about 110 Da then molecular weight of protein with 160 amino acids = 160 x 110 = 17,600 D = 17600/1000 KDa =17.6 KDa